# Searching for Particle Dark Matter with eROSITA Early Data

### v2

@@ Abstract
Many well motivated dark matter (DM) particle candidates can decay into detectable X-ray photons. We analyze the Final Equatorial Depth Survey (eFEDS) data from eROSITA early data release to search for unidentified X-ray lines that could indicate DM signals. Having discovered no anomalous signal, we set limits on DM decay rate in mass range between 1.8-18 keV, and constrain the parameter space of two DM particles: sterile neutrino and axion-like particles. Finally we also study the projected sensitivity of eROSITA full sky search, showing that eROSITA all-sky survey is expected to set the most stringent limits in the soft X-ray band.
@@ I Introduction
X-ray searches of DM had gained significant attention in the past decade, when a stacked X-ray analysis of XMM-Newton studying 73 galactic clusters indicated there’s an unexplained line around , which could correspond to the decay signature of a DM particle with mass at Bulbul et al. (2014). line was also claimed in 3 later studies: XMM-Newton observation of Andromeda and Perseus cluster Boyarsky et al. (2014), XMM-Newton observation of Milky way center Boyarsky et al. (2015), and Chandra deep field observation Cappelluti et al. (2018). However, other studies could not find line or any other indication of DM decay Horiuchi et al. (2014); Urban et al. (2015); Malyshev et al. (2014); Anderson et al. (2015); Tamura et al. (2015); Jeltema and Profumo (2015, 2016); Aharonian et al. (2017); Gewering-Peine et al. (2017); Silich et al. (2021); Sicilian et al. (2020, 2022); Dessert et al. (2020); Foster et al. (2021); Roach et al. (2023). Recently, a study Dessert et al. (2023) attempted to reproduce the 3 original analyses where was discovered, and could not found robust evidence to support its existence. Nevertheless, x-ray line searches remains a important way of searching for DM signatures.
@@
In this paper, we perform indirect DM search on the largest continuous observation from publicly available eROSITA data. We search for DM decay signal in eROSITA Equatorial Field Depth Survey (eFEDS) data. In the event of null detection, we set upper limit on the decay rate of DM. We also convert the decay rate limit to parameter space constraints for sterile neutrino mixing angle and Axion-Like Particles couplings in mass range -. Finally, we extrapolate the result that would be obtained from eROSITA using 4 years of full sky data. This study also serves as a preparation for DM analysis on eROSTIA full data release.
@@
eROSITA consists of 7 independent telescope modules (TM1-TM7), arranged in a honey comb like geometry Eder et al. (2018). A TM contains a mirror assembly and a camera module. The mirror assembly has 54 paraboloid and hyperboloid mirror shells in a Wolter-I geometry. There’s a camera at the focus of the mirror. It has a circular field of view with diameter at 1.03 degrees, which makes the geometric area of the detector 0.833 deg. The grasp of eROSITA, defined as the energy dependent multiplied product of effective area and field of view, is competitive in 3.5-10 keV range, while surpassing other instruments in 0.3-3.5 keV Predehl et al. (2021).
@@
For this analysis, we ignore the energy channels below 0.9 keV in our spectral fitting and unexplained X-ray DM line analysis. This solution is straightforward but effective. In doing so, we make use of data taken by all 7 modules, while mitigate light leak effect such that our blank sky spectrum is still valid.
@@
eROSITA obsID |
Central RA |
Central DEC |
Effective Exposure111during this time all 7 TMs were working |
300007-1 |
129.5500 |
1.5000 |
89642.0 |
300008-1 |
133.8600 |
1.5000 |
89642.0 |
300009-1 |
138.1400 |
1.5000 |
89642.0 |
300010-1 |
142.4500 |
1.5000 |
89642.0 |
@@
We first use evtool from eSASS to combine the 4 observations of eFEDS into one stacked observation, with setting: flag=0xC000F000 to remove unusually bright pixels, gtitype=flaregti to remove time intervals with flares, and pattern=15 for photon reconstruction pattern (following the procedure in appendix A of Brunner et al. (2022)). We expect stacking these 4 observations would not influence the accuracy of sky background or DM signal model, since their locations are within a few degrees of each other. Using srctool, we then obtain 7 spectra with unit in ph/kev/s, one for each TM, as well as their corresponding Response Matrix Function (RMF) and Auxiliary Response Function (ARF) files. This allows us to account for 7 TM’s slightly different internal background and response matrix. For srctool, we set parameter exttype=point. The target region is set to a box sized 20 by 10, centered at eFEDS location, to encompass the entirety of eFEDS field. The details of stacked eFEDS observation after processing are summarized in table 2.
Telescope Module |
Central RA |
Central DEC |
Combined, Cleaned Effective Exposure |
Combined, Cleaned Sky area |
TM1 |
136.00 |
1.50 |
360312 |
164.46 |
TM2 |
136.00 |
1.50 |
360583 |
163.52 |
TM3 |
136.00 |
1.50 |
357331 |
164.47 |
TM4 |
136.00 |
1.50 |
350152 |
164.49 |
TM5 |
136.00 |
1.50 |
301490 |
164.44 |
TM6 |
136.00 |
1.50 |
293827 |
164.42 |
TM7 |
136.00 |
1.50 |
360543 |
164.46 |
@@
The final step of our data reduction is rebinning. Each spectra is rebinned to equal log width bins, specifically 200 bins per decade. The binning scheme is chosen such that the bins are narrower than one photon energy resolution (FWHM), but also large than the many-photon energy resolution (FWHM/). For eROSTIA, FWHM ranges from eV at 1 keV, to eV at 8 keV. After binning for individual TMs, each bin in our spectrum has roughly between 2000 to 6000 photons.
@@ II.3 Modeling eFEDS Blank Sky
We perform the spectral fitting with XSPEC, version number 12.12.1 Arnaud (1996). Our null sky model has two parts, instrumental background and astrophysical background (also named Cosmic X-ray Background (CXB) following Liu et al. (2022)).
The instrumental background is the dominant component above keV. It is caused by various cosmic particles interacting with the instrument as well as other noises. The model is based on in-flight measurement taken during Cal-PV period, with camera filter wheel set to “closed” position. Both the data and model of filter wheel closed observation are produced by eROSITA collaboration, and are made publicaly available on EDR website 444https://erosita.mpe.mpg.de/edr/eROSITAObservations/EDRFWC/. The model has four components to describe the continuum, a double broken power law and two additional power laws, all multiplied by an exponential modification. On top of the continuum there are 14 Gaussian lines to model emission from various internal elements Whelan et al. (2022). During spectral fitting, we fix the three power law indices, the exponential modification factor, and two break energies describing the continuum, following the precedent set by previous eROSITA work Whelan et al. (2022). All 4 normalization parameters for continuum are allowed to float for 5 TMs unaffected by light leak. For TM5 and TM7 there’s a power law normalization frozen at 0, following FWC model provided by eROSITA EDR. For the instrument Gaussian lines, we keep the line energy location and line width fixed, only allowing the normalization to float. We assume no correlation between free model parameters across different TMs and allow them to fit independently. See appendix A for a full table of frozen parameters and their values.
@@
For astrophysical X-ray background, we consider (1) an absorbed apec Smith et al. (2001) model with temperature for the hot galactic halo emission, and (2) an absorbed power law with index fixed at 1.46 as cosmic X-ray background accounting for unresolved sources. This modeling choice is based on previous eROSITA eFEDS analysis Liu et al. (2022), which itself is based on XMM-Newton CXB model in soft X-ray range Kuntz and Snowden (2000); Snowden et al. (2008); Bulbul et al. (2012). There is a component in the previous models that we do not include, which is an unabsorbed apec model with keV for the thermal emission from local hot bubble. Since we ignore data keV, including this extra apec in the model will only change the best fit by . In the two CXB components, all normalization parameters are free to vary. The temperature of the hot galactic halo apec model is allowed to vary within ranges of 0.5-1 keV. We use model tbabs Wilms et al. (2000) for intergalactic medium absorption, with hydrogen column density value at eFEDS target location fixed at 4.4e20 cm and abundance fixed at solar, adopting abundance table from Wilms et al. (2000). The calculation of hydrogen column density value is provided by UK Swift Science Data Centre Willingale et al. (2013). In total our blank sky model has 20 free parameters. The free parameters of the blank sky model and their ranges are summarized in table 3.
@@
Source | Component name | XSPEC model name | Free Parameters |
CXB | absorbed555Absorption from interstellar medium, modeled by tbabs. Same for the power law unresolved point sources below. hotter galactic halo | apec | Plasma temperature , keV |
(0.5-1) | |||
Normalization | |||
(Free non-negative value) | |||
CXB | absorbed power law, from un- | powerlaw | Normalization |
resolved, diffuse AGN sources | (Free non-negative value)) | ||
Instrument | Internal Continuum | expfac(bkn2pow + powerlaw + powerlaw) | Normalization |
(Free non-negative value) | |||
Instrument | Internal Gaussian lines | gaussian | Normalization |
(Free non-negative value) |
@@ III Dark Matter Analysis
To search for a possible signature in eFEDS from DM decay, we fit the spectrum with a mock signal on top of our null model. The mock signal has a Gaussian shape, with normalization determined by free parameter decay rate and fixed parameter , see equations in section III.2. Since we assume the origin of the line is from galactic DM, the Gaussian line width is set at 0, consistent with energy resolution of eROSITA, for Predehl et al. (2021). The green and grey lines in figure 2 demonstrate of our fiducial signal, and a refitted model with said signal.
@@
The line search procedure shows eFEDS observation is consistent with null detection model. We then turn to interpret this result in the context of galactic DM search, and set upper limit on DM decay rate.
@@ III.2 Dark Matter Signal Modeling
where is the DM density profile, is galactocentric radius, is line of sight distance, is pointing angle from galactic center, and kpc is the distance of solar system to Galactic center.
@@
where kpc is scaling distance Hooper (2017) and GeV/cm is DM density at solar neighborhood de Salas et al. (2019); de Salas and Widmark (2021); Sofue (2020). For Milky Way kinetic simulation, the original best fit parameter shows Navarro et al. (1997). We denote this DM density profile with a cusp at the center as simply NFW. Numerical simulation with baryon added in shows the inner slope of DM could have a core Calore et al. (2015); Nesti and Salucci (2013). To approximate this case, we use an NFW profile where the slope is shallow all the way to the center with Pato et al. (2015) denoted as shallow NFW. We also consider Einasto profile Diemand et al. (2008) described by
with kpc. The typical Einasto profile (EIN) has . Recently, two studies calculated best fit DM density profile from Gaia DR3 Lim et al. (2023); Ou et al. (2023). However, the best fit DM density and D-factor obtained from these studies are noticeably lower than previous results, as shown in figure 3. To ensure our DM analysis is consistent with previous DM studies and can be compared, we do not make use of these latest DM density profiles from Gaia DR3. For our analysis we take NFW profile as the most general description for DM and only report one sets of results. We note that depends on the specific DM density model, our constraint result could be modified by a factor of 2.
@@ III.3 Dark Matter Decay Rate Constraint and Sensitivity Analysis
We obtain DM decay rate upper limit by using the same scanning and fitting procedure described in III.1. At each of the 200 steps, we scan through DM line normalization value, start with and increasing it until the best fit model with DM line has . corresponds to the 95% one sided upper limit on DM decay flux.
@@
Overall, most regions of our limit falls within 2 containment region. There are a few places where it goes beyond. In particular, the limit exceeds 2 containment at around -. At this energy range, we argue it reflects a modeling shortcomings in spectral energy -. As shown in figure 9 and 10 of Predehl et al. (2021), the effective area, as well as grasp, of eROSITA detector drop off sharply between -, followed by making a tiny bump. The quirks in our limit is a direct consequence of effective area calibration done by eROSITA collaboration. One recent eROSITA study on 3.5 keV line pointed out this feature could come from mismodeling of gold absorption edge Villalba (2024). With an improved effective area model, we expect this deviation to disappear in future analysis. In the current work, we choose to not report the limit if it goes below containment, shown on figure 4 as grey vertical bands.
III.4 All Sky Projection with Mock Data
We show the DM decay rate sensitivity projection of eROSITA all-sky and galactic center survey in figure 5. Using XSPEC fakeit function, we generate a set of mock spectra based on our best fit models, assuming no fluctuation, and obtain limit from it. This represents decay rate limit if we can perfectly model eFEDS spectra. Assuming null detection hypothesis holds for all regions of the sky, the decay lifetime limit has the simple relation , since eFEDS is designed to have the same observation depth as the final all sky survey Brunner et al. (2022).
Next, we generate a sky map using the software package healpy Górski et al. (2005); Zonca et al. (2019). With parameter Nside = 5, the generated sky map that has 300 equal area pixels, each having sky area. We set the closest pixel to Galactic center at 4 away from , to avoid NFW profile’s DM density cusp as well as the potential strong astrophysical emission in the actual data. Next we create a projected at each pixel based on D-factor at its coordinate and size. Finally we and find where . We will discuss our projection in the context of sterile neutrino and ALP DM searches in later sections.

### v3

@@ Abstract
Many well motivated dark matter (DM) particle candidates can decay into detectable X-ray photons. We analyze the Final Equatorial Depth Survey (eFEDS) data from eROSITA early data release to search for unidentified X-ray lines that could indicate DM signals. Having discovered no anomalous signal, we set limits on DM decay rate in mass range between 1.8-18 keV, and constrain the parameter space of two DM particles: sterile neutrino and Axion-like particle. Finally we also study the projected sensitivity of eROSITA full sky search, showing that eROSITA all-sky survey is expected to set the most stringent limits in the soft X-ray band.
@@ I Introduction
X-ray searches of DM had gained significant attention in the past decade, when a stacked X-ray analysis of XMM-Newton studying 73 galactic clusters indicated there’s an unexplained line around , which could correspond to the decay signature of a DM particle with mass at Bulbul et al. (2014). line was also claimed in 3 later studies: XMM-Newton observation of Andromeda and Perseus cluster Boyarsky et al. (2014), XMM-Newton observation of Milky Way Center Boyarsky et al. (2015), and Chandra deep field observation Cappelluti et al. (2018). However, other studies could not find line or any other indication of DM decay Horiuchi et al. (2014); Urban et al. (2015); Malyshev et al. (2014); Anderson et al. (2015); Tamura et al. (2015); Jeltema and Profumo (2015, 2016); Aharonian et al. (2017); Gewering-Peine et al. (2017); Silich et al. (2021); Sicilian et al. (2020, 2022); Dessert et al. (2020); Foster et al. (2021); Roach et al. (2023). Recently, a study Dessert et al. (2023) attempted to reproduce the 3 original analyses where was discovered, and could not NFWfound robust evidence to support its existence. Nevertheless, X-ray line searches remain a important way of searching for DM signatures.
@@
In this paper, we perform indirect DM search on the largest continuous observation from publicly available eROSITA data. We search for DM decay signal in eROSITA Equatorial Field Depth Survey (eFEDS) data. In the event of null detection, we set upper limit on the decay rate of DM. We also convert the decay rate limit to parameter space constraints for sterile neutrino mixing angle and Axion-Like Particles couplings in mass range -. Finally, we extrapolate the result that would be obtained from eROSITA using 4 years of full sky data. This study also serves as a preparation for DM analysis on eROSITA full data release.
@@
eROSITA consists of 7 independent telescope modules (TM1-TM7), arranged in a honey comb like geometry Eder et al. (2018). A TM contains a mirror assembly and a camera module. The mirror assembly has 54 paraboloid and hyperboloid mirror shells in a Wolter-I geometry. There’s a camera at the focus of the mirror. It has a circular field of view with diameter at 1.03 degrees, which makes the geometric area of the detector 0.833 deg2. The grasp of eROSITA, defined as the energy dependent multiplied product of effective area and field of view, is competitive in 3.5-10 keV range, while surpassing other instruments in 0.3-3.5 keV Predehl et al. (2021).
@@
For this work, we ignore the energy channels below 0.9 keV in our spectral fitting procedure and in our DM analysis. This solution is straightforward but effective. In doing so, we make use of data taken by all 7 modules, while mitigate light leak effect such that our blank sky spectrum is still valid.
@@
| eROSITA obsID | Central RA | Central DEC | Effective Exposure111during this time all 7 TMs were working |
| (deg) | (deg) | (s) | |
| 300007-1 | 129.5500 | 1.5000 | 89642.0 |
| 300008-1 | 133.8600 | 1.5000 | 89642.0 |
| 300009-1 | 138.1400 | 1.5000 | 89642.0 |
| 300010-1 | 142.4500 | 1.5000 | 89642.0 |
@@
We first use evtool from eSASS to combine the 4 observations of eFEDS into one stacked observation, with setting: flag=0xC000F000 to remove unusually bright pixels, gtitype=flaregti to remove time intervals with flares, and pattern=15 for photon reconstruction pattern (following the procedure in appendix A of Brunner et al. (2022)). We expect stacking these 4 observations would not influence the accuracy of sky background or DM signal model, since their locations are within a few degrees of each other. Using srctool, we then obtain 7 spectra with unit in ph/kev/s, one for each TM, as well as their corresponding Response Matrix Function (RMF) and Auxiliary Response Function (ARF) files. This allows us to account for 7 TM’s slightly different internal background and response matrix. For srctool, we set parameter exttype=point. The target region is set to a box sized 20∘ by 10∘, centered at eFEDS location, to encompass the entirety of eFEDS field. The details of stacked eFEDS observation after processing are summarized in table 2.
| Telescope Module | Central RA | Central DEC | Combined, Cleaned Effective Exposure | Combined, Cleaned Sky area |
| (deg) | (deg) | (s) | (deg2) | |
| TM1 | 136.00 | 1.50 | 360312 | 164.46 |
| TM2 | 136.00 | 1.50 | 360583 | 163.52 |
| TM3 | 136.00 | 1.50 | 357331 | 164.47 |
| TM4 | 136.00 | 1.50 | 350152 | 164.49 |
| TM5 | 136.00 | 1.50 | 301490 | 164.44 |
| TM6 | 136.00 | 1.50 | 293827 | 164.42 |
| TM7 | 136.00 | 1.50 | 360543 | 164.46 |
@@
The final step of our data reduction is rebinning. Each spectra is rebinned to equal log width bins, specifically 200 bins per decade. In our binning scheme, the bin width is smaller than eROSITA energy resolution (FWHM), so that a Gaussian line would show up in multiple bins. For eROSITA, FWHM ranges from eV at 1 keV, to eV at 8 keV. After binning for individual TMs, each bin in our spectrum has roughly between 2000 to 6000 photons. We have tried using the default binning scheme given by eROSITA data, which is used by RMF and has 728 bins between 0.9 - 9 keV. In this binning scheme, there are 700 to 2500 photon counts in each bin. Between these two choices of binning, we see no significant difference in the best fit /d.o.f. and DM decay rate constraint.
@@ II.3 Modeling eFEDS Blank Sky
We perform the spectral fitting with XSPEC, version number 12.12.1 Arnaud (1996). Our null sky model has two parts, instrumental background and astrophysical background.
The instrumental background is the dominant component above keV. It is caused by various cosmic particles interacting with the instrument as well as other noises. The model is based on in-flight measurement taken during Cal-PV period, with camera filter wheel set to “closed” position. Both the data and model of filter wheel closed observation are produced by eROSITA collaboration, and are made publicly available on EDR website 444https://erosita.mpe.mpg.de/edr/eROSITAObservations/EDRFWC/. The model has four components to describe the continuum, a double broken power law and two additional power laws, all multiplied by an exponential modification. On top of the continuum there are 14 Gaussian lines to model emission from various internal elements Whelan et al. (2022). During spectral fitting, we allow the three power law indices, the exponential modification factor, two break energies, and all 4 normalization parameters for continuum to float freely. The only exception is the index and normalization of the last power law for TM5 and TM7, which are frozen at 6 and 0, following FWC model provided by eROSITA EDR. For the instrument Gaussian lines, we fix the line widths, allow the line energy location to float within ranges preset by eROSITA FWC model, and allow the normalizations to float freely. We assume no correlation between free model parameters across different TMs and allow them to fit independently.
@@
For astrophysical X-ray background, we consider (1) an absorbed apec Smith et al. (2001) model with temperature for the Local Group emission, and (2) an absorbed power law with index around 1.46 accounting for unresolved extragalactic sources, sometimes called Cosmic X-ray background (CXB). This modeling choice is based on previous eROSITA eFEDS analysis Liu et al. (2022), which itself is based on XMM-Newton model in soft X-ray range Kuntz and Snowden (2000); Snowden et al. (2008); Bulbul et al. (2012). In the two astrophysical components, all normalization parameters are free to vary. The temperature of the hot galactic halo apec model is allowed to vary within range of 0.5-1 keV. We use model tbabs Wilms et al. (2000) for intergalactic medium absorption, with he initial fit value of hydrogen column density value at eFEDS target location set to 4.4e20 cm-2 and initial fit value of abundance set to solar, adopting abundance table from Wilms et al. (2000). The calculation of hydrogen column density value is provided by UK Swift Science Data Centre Willingale et al. (2013). To account for uncertainty from modeling, we allow all parameters to float, and assume there is no correlation between the parameters for each TM. The best fit parameters of our blank sky model roughly agree with the numbers reported in eROSITA diffuse sky study Ponti et al. (2023). See the appendix table 3 for the exact values of our best fit parameters and their allowed ranges. We also test the impact of fixing CXB power law index to values reported in previous literature. We find that, the choices of different CXB power law indices change our DM constraint result by less than factor of 3 in most energies. See appendix B and figure B.1 within for further discussion.
There is another unabsorbed apec model component commonly present in previous astrophysical X-ray background modeling, meant to model Local Hot Bubble Snowden et al. (2008). This apec peaks at around - keV, which is well below the - keV energy range where we fit the spectrum and search for DM. We have tried fitting eFEDS spectrum with this unabsorbed apec component included alongside the absorbed apec and power law. The best fit and the DM search constraint show no noticeable change except near the low energy edge. Therefore, we conclude it is safe to exclude this component in our blank sky modeling.
@@ III Dark Matter Analysis
To search for a possible signature in eFEDS from DM decay, we fit the spectrum with a mock signal on top of our null model. The mock signal has a Gaussian shape, with normalization determined by free parameter decay rate and fixed parameter , see equations in section III.2. Regarding DM Gaussian line width, it is important to consider the potential Doppler shift resulted from DM velocity dispersion in Milky Way halo. Recent works find that the Doppler shift from velocity dispersion will result in DM line broadening Speckhard et al. (2016); Powell et al. (2017); Zhong et al. (2020); Dessert et al. (2024); Lovell (2024). This level of broadening is far lower than eROSITA energy resolution, which is between 2%-8% in the energy range we study Predehl et al. (2021). Therefore, in our analysis, we did not consider the potential line broadening from DM dispersion velocity, and set DM Gaussian line width to be 0 in our line search procedure. With DM Gaussian line parameters determined, we then fold the model through corresponding instrument responses functions (RMF and ARF) extracted with srctool, similar to all other models with astrophysical origin. The green and grey lines in figure 2 demonstrate of our fiducial signal, and a refitted model with said signal.
@@
The line search procedure shows eFEDS observation is mostly consistent with our null model, with a notable exception of an excess at spectral energy 1.9 keV. Due to this excess, the best fit with DM indicates for at decay rate . We do not believe this feature indicates a detection of DM or unexplained X-ray line, but rather mis-modeling of eROSITA’s rapidly changing effective area in this energy range, also reported in other eFEDS studies (e.g., figure 6 of Liu et al. (2022)). eFEDS study on cluster Abell 3266 reported 10% residual and suspected it’s due to calibration uncertainty of the telescope Sanders et al. (2022). We will discuss the impact of these unexpected features on our DM constraint and our treatment in section III.3.
Next we turn to interpret the line search result in the context of galactic DM decay, and set upper limit on DM decay rate.
@@ III.2 Dark Matter Signal Modeling
where is the DM density profile, is galactocentric radius, is line of sight distance, is pointing angle from Galactic Center, and kpc is the distance of solar system to Galactic Center. In our DM analysis, we evaluated the integral in eq. 1 by dividing the eFEDS observation region using healpy Górski et al. (2005); Zonca et al. (2019) with nside parameter = 128. This gives us 776 sub regions, with each being 0.21 deg2. Then for each healpy pixel we calculated the D factor and summed over them.
@@
where kpc is scaling distance Hooper (2017) and GeV/cm3 is DM density at solar neighborhood de Salas et al. (2019); de Salas and Widmark (2021); Sofue (2020). For Milky Way kinetic simulation, the original best fit parameter shows Navarro et al. (1997). We denote this DM density profile with a cusp at the center as simply NFW. Numerical simulation with baryon added in shows the inner slope of DM could have a core Calore et al. (2015); Nesti and Salucci (2013). To approximate this case, we use an NFW profile where the slope is shallow all the way to the center with Pato et al. (2015) denoted as shallow NFW.
We also consider Einasto profile Diemand et al. (2008) described by
with kpc. The typical Einasto profile (EIN) has .
Lastly, we consider Burkert profile which was first introduced to model dwarf galaxy Burkert (1995)
where is the characteristic density and is the scale length. We take the best fit parameters and estimated by a study using LAMOST and SDSS observations Lin and Li (2019).
Recently, two studies calculated best fit DM density profile from Gaia DR3 Lim et al. (2023); Ou et al. (2023). However, the best fit DM density and D-factor obtained from these studies are noticeably lower than previous results, as shown in figure 3. To ensure our DM analysis is consistent with previous DM studies and can be compared, we do not make use of these latest DM density profiles from Gaia DR3. For our analysis we take NFW profile as the most general description for DM and only report one sets of results. We note that depends on the specific DM density model, our constraint result could be modified by a factor of 2.
@@ III.3 Dark Matter Decay Rate Constraint and Sensitivity Analysis
We obtain DM decay rate upper limit by using the same scanning and fitting procedure described in III.1. At each of the 200 steps, we scan through DM line normalization value, start with and increasing it until the best fit model with DM line has
In this equation, we look for , which corresponds to the 95% one sided upper limit on DM decay flux.
@@
Overall, most regions of our limit falls within 2 containment region. There are a few places where it goes beyond. As shown in the Monte Carlo analysis, our limit exceeds two sigma containment at 4-5 keV. The energy in photon spectrum corresponding to the problematic DM mass is 2-2.5 keV, which shows two “troughs” and a “bumps”. We think all of them are due to inaccuracies in calibration for effective area by eROSITA collaboration at this energy range, as discussed in section III.1. In particular, the limit exceeds 2 containment at around -. At this energy range, we argue it reflects a modeling shortcomings in spectral energy -. As shown in figure 9 and 10 of Predehl et al. (2021), the effective area, as well as grasp, of eROSITA detector drop off sharply between -, followed by making a tiny bump. The quirks in our limit is a direct consequence of effective area calibration done by eROSITA collaboration. One recent eROSITA study on 3.5 keV line investigated the possibility of this feature coming from mis-modeling of gold absorption edge, but ultimately concluded gold edge alone is not enough to explain the residual Villalba (2024). With an improved effective area model, we expect this deviation to disappear in future analysis.
Our procedure for line search and setting constraints would deviate from Monte Carlo expectation when the observed photon spectrum has deficit and excess from the model. When the observed photon spectrum has an excess over the model, our limits become weaker than expected, and thus is conservative, but our line search procedure still can rule out DM signal above the constraint. As it would not falsely rule out the presence of a signal due to the profiling procedure in our analysis, we keep the conservative limit in our results. On the other hand, since the deficits would cause the constraint to mistakenly rule out DM parameters, we choose to exclude the overly stringent limit below containment. This regions are marked on figure 4 as grey vertical bands.
III.4 Future eROSITA Survey Projection with Mock Data
Figure 5 shows our projected future eROSITA survey sensitivity to DM decay based our result from eFEDS. We also show the limit obtained from eFEDS itself and from mock data on figure 5 for comparison.
We obtain the projected sensitivity using mock spectra generated for each pixel on a sky map. Specifically, we generate a sky map with 300 pixels, using healpy Górski et al. (2005); Zonca et al. (2019) with parameter Nside=5. Each pixel is 126 deg2 large, roughly the same order of magnitude as eFEDS observation at 163 deg2. We exclude the center pixel in this sky map situated at for our analysis, since in this region DM density profile has the largest uncertainty.
For projected Galactic Center (GC) sensitivity, we use the pixel on the healpy sky map that does not cover , instead situated at making its galactocentric angle . We generate 7 mock spectra using the XSPEC’s fakeit function, corresponding to 7 TMs of eROSITA. We assume eROSITA will observe this region with the same exposure as eFEDS at 360 ks. The mock spectra follow the same best fit model and parameters for each TM used in our main eFEDS analysis, and assumes the same systematic error. We pass the mock spectra through the same line search and DM analysis procedure as the main eFEDS analysis described in section III.3. The obtained limit is our projected GC sensitivity, with the D-factor calculated using the pixel’s center point.
For all-sky survey sensitivity projection, we take the sky map with 299 pixels (after excluding the center pixel), as the total sky area observed by eROSITA suitable for DM analysis. Similar to the GC survey pixel, we assume in the future eROSITA will complete all-sky survey by observing every pixel on the map with the same exposure time as eFEDS at 360 ks. Next, separately for each pixel, we generate 7 mock spectra using the XSPEC’s fakeit function, making a total of 2093 mock spectra representing all the data from eFEDS’s all-sky survey. Finally, we pass these mock spectra through a slightly modified DM decay rate constraint procedure. For all-sky sensitivity, we sum over pixel index, and find the projected limit by:
In this equation, corresponds to the summed best fit from 7 TM mock data, and corresponds to the pixel index going from 1 to 299. The profile is calculated individually for each pixel with the D-factor calculated using the center point of the pixel.
We note that our procedure, for both GC and all-sky projection, ignores the background effect coming from Galactic disk and point sources. For DM line search performed in this work, we expect the impact of continuum background component to be relatively minor. Line background on the other hand, could be important due to signal background confusion. Thus the projected sensitivities reported in figure 5 represent slightly optimistic cases. We defer a detailed calculation taken all background component into account to future eROSITA DM analysis. Taken at face value, eROSITA’s GC sensitivity improves from eFEDS observation by 10, which is the ratio of D-factor. The all-sky sensitivity on the other hand roughly improves from eFEDS by ratio of all-sky average D-factor times , as a result of increase in statistics. We will discuss the specific implications of these projections in the next sections.